OpenStudy (anonymous):

(3/xy^2)^-3 simplify?

OpenStudy (bohotness):

Step by step solution : Step 1 : Raise (2xy) to the 3rd power Exponentiaion : Equation at the end of step 1 : ((3x • (y2)) • 4xy) • 23x3y3 Step 2 : Raise y to the 2nd power Exponentiaion : Equation at the end of step 2 : ((3x • y2) • 4xy) • 23x3y3 Step 3 : Multiply 3xy2 by 4xy Multiplying exponential expressions : 3.1 x1 multiplied by x1 = x(1 + 1) = x2 Multiplying exponential expressions : 3.2 y2 multiplied by y1 = y(2 + 1) = y3 Equation at the end of step 3 : (3•22x2y3) • 23x3y3 Step 4 : Multiply (3•22x2y3) by 23x3y3 Multiplying exponential expressions : 4.1 x2 multiplied by x3 = x(2 + 3) = x5 Multiplying exponential expressions : 4.2 y3 multiplied by y3 = y(3 + 3) = y6 Multiplying exponents : 4.3 22 multiplied by 23 = 2(2 + 3) = 25